# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix D Proof of Lipschitz-continuity Bound
We extend $a_{\Delta}^{*}$ to $a_{\Delta^{\prime}}^{*}$ by letting $a_{\Delta^{\prime}\setminus\Delta}^{*}$ be the corresponding $\Delta\setminus\Delta^{\prime}$ variables from $\widetilde{a}_{\Delta^{\prime}}^{*}$ in $\mathcal{A}_{l}^{|\Delta^{\prime}\setminus\Delta|}$. For the remainder of this proof, we adopt the shorthand $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}$ to refer to $\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}$. Then, if $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})>0$, we have:
$\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})$ $-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})\bigg{|}$
$=\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{*},a^{*}_{\Delta})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},\widetilde{a}_{g}^{*},\widetilde{a}^{*}_{\Delta^{\prime}})$
$\leq\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}^{T}_{k}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{*},a_{\Delta}^{*})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{*},a_{\Delta^{\prime}}^{*})$
$\leq\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{|}$
@@
Similarly, if $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})<0$, an analogous argument by replacing $a_{g}^{*}$ with $\widetilde{a}_{g}^{*}$ and $a_{\Delta}^{*}$ with $\widetilde{a}_{\Delta}^{*}$ yields an identical bound. ∎
Lemma D.13.
@@
Let $\widehat{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})$ and $\widetilde{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})$.
@@
$=\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\widehat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\widetilde{\theta}^{*})$
$\leq\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\widehat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\widehat{\theta}^{*})$
@@
Here, we replace each $\theta^{*}$ with the maximizers of their corresponding terms, and upper bound them by the maximizer of the larger term. Next, we replace $\widehat{\theta^{*}}$ in both expressions with the maximizer choice $\theta^{*}$ from $\Theta^{*}$, and further bound the expression by its absolute value.
If $\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}},\dots,\theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}},\dots,\theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})<0$, then an analogous argument that replaces $\widehat{\theta}^{*}$ with $\widetilde{\theta}^{*}$ yields the same result. ∎
Appendix E Bounding Total Variation Distance
As $|\Delta|\to n$, we prove that the total variation (TV) distance between the empirical distribution of $z_{\Delta}$ and $z_{[n]}$ goes to 00. Here, recall that $z_{i}\in\mathcal{Z}=\mathcal{S}_{l}\times\mathcal{A}_{l}$, and $z_{\Delta}=\{z_{i}:i\in\Delta\}$ for $\Delta\in{[n]\choose k}$. Before bounding the total variation distance between $F_{z_{\Delta}}$ and $F_{z_{\Delta^{\prime}}}$, we first introduce Lemma C.5 of Anand & Qu (2024) which can be viewed as a generalization of the Dvoretzky-Kiefer-Wolfowitz concentration inequality, for sampling without replacement. We first make an important remark.
Remark E.1.
@@ Appendix E Bounding Total Variation Distance
For reference, we present the multidimensional Dvoretzky-Kiefer-Wolfowitz (DKW) inequality (Dvoretzky et al. (1956); Massart (1990); Naaman (2021)) which bounds the difference between an empirical distribution function for a set $B_{\Delta}$ and $B_{[n]}$ when each element of $\Delta$ for $|\Delta|=k$ is sampled uniformly at random from $[n]$ with replacement.
Theorem E.2 (Multi-dimensional Dvoretzky-Kiefer-Wolfowitz (DFW) inequality (Dvoretzky et al., 1956; Naaman, 2021)).
@@
Lemma C.5 of Anand & Qu (2024) generalizes the DKW inequality for sampling without replacement:
Lemma E.3 (Sampling without replacement analogue of the DKW inequality, Lemma C.5 in Anand & Qu (2024)).
@@
Lemma E.4.
With probability atleast $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$, $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})\leq\epsilon$.
@@
Recall that $z:=(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$. From Lemma E.3, substituting $\mathcal{B}_{l}=\mathcal{Z}_{l}$ yields:
@@
Lemma E.5 (Lemma B.8.1 of Ghosal & van der Vaart (2017). Subadditivity of TV distance for Product Distributions).
@@
Furthermore, we will need the following lemma.
Lemma E.6 (Data Processing Inequality.).
Let $A$ and $B$ be random variables over some domain $S$. Let $f$ be some function (not necessarily deterministically) mapping from $S$ to any codomain $T$. Then,
$\mathrm{TV}(f(A),f(B))\leq\mathrm{TV}(A,B)$
Lemma E.7.
@@
From the Bretagnolle-Huber inequality Tsybakov (2008) we have that
@@
In the third line, we note that $\sum_{z\in\mathcal{Z}_{l}}\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}=|\Delta|$ since each local agent contained in $\Delta$ must have some state/action pair contained in $\mathcal{Z}_{l}$. In the last line, we note that $\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\leq\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}$, For all $z\in\mathcal{Z}_{l}$, and thus the summation of logarithmic terms in the third line is negative. Finally, using this bound in the Bretagnolle-Huber inequality yields the lemma.∎
Corollary E.7.1.
From Lemma E.7, setting $\Delta=[n]$ also recovers $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})=0$.
Theorem E.8.
@@
$\left|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
@@
From combining the total variation distance bound in Lemma E.4 and the Lipschitz continuity bound in Theorem D.3 with $\sum_{t=0}^{T}\gamma^{T}\leq\frac{1}{1-\gamma}$ for $\gamma\in(0,1)$, we have:
$\Pr\left[|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})|\leq\frac{2\epsilon}{1-\gamma}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}\right]\geq 1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$
We then reparameterize $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$ into $1-\delta$ to get $\epsilon=\sqrt{\frac{n-k+1}{8kn}\ln\left(\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}\right)}$ which gives:
$\Pr\left[\left|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}\right]\geq 1-\delta,$
@@
Appendix F Using the Performance Difference Lemma
In general, convergence analysis requires the guarantee that a stationary optimal policy exists. Fortunately, when working with the empirical distribution function, the existence of a stationary optimal policy is guaranteed when the state/action spaces are finite or countably infinite. However, lifting the knowledge of states onto the continuous empirical distribution function space, and designing a policy on the lifted space is still analytically challenging. To circumvent this, Gu et al. (2021) creates lifted $\epsilon$-nets and does kernel regression to obtain convergence guarantees. Here, our analytic approach bears a stark difference, wherein we analyze the sampled structure of the mean-field empirical distribution function, rather than studying the structure of the lifted space. For this, we leverage the classical Performance Difference Lemma, which we restate below for completeness.
Lemma F.1 (Performance Difference Lemma, Kakade & Langford (2002)).

### v3

@@ Appendix E Proof of Lipschitz-continuity Bound
We extend $a_{\Delta}^{*}$ to $a_{\Delta^{\prime}}^{*}$ by letting $a_{\Delta^{\prime}\setminus\Delta}^{*}$ be the corresponding $\Delta\setminus\Delta^{\prime}$ variables from $\tilde{a}_{\Delta^{\prime}}^{*}$ in $\mathcal{A}_{l}^{|\Delta^{\prime}\setminus\Delta|}$. For the remainder of this proof, we adopt the shorthand $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}$ to refer to $\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}$. Then, if $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})>0$, we have:
$\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})\bigg{|}$
$=\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{*},a^{*}_{\Delta})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},\tilde{a}_{g}^{*},\tilde{a}^{*}_{\Delta^{\prime}})$
$\leq\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}^{T}_{k}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{*},a_{\Delta}^{*})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{*},a_{\Delta^{\prime}}^{*})$
$\leq\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{|}$
@@
We observe that if the opposite inequality holds (i.e., $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})<0$), an analogous argument by replacing $a_{g}^{*}$ with $\tilde{a}_{g}^{*}$ and $a_{\Delta}^{*}$ with $\tilde{a}_{\Delta}^{*}$ yields an identical bound. ∎
Lemma E.15.
@@
Let $\hat{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})$ and $\tilde{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})$.
@@
$=\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\hat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\tilde{\theta}^{*})$
$\leq\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\hat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\hat{\theta}^{*})$
@@
Here, we replace each $\theta^{*}$ with the maximizers of their corresponding terms, and upper bound them by the maximizer of the larger term. Next, we replace $\hat{\theta^{*}}$ in both expressions with the maximizer choice $\theta^{*}$ from $\Theta^{*}$, and further bound the expression by its absolute value.
If $\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}},\dots,\theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}},\dots,\theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})$ is negative, then an analogous argument that replaces $\hat{\theta}^{*}$ with $\tilde{\theta}^{*}$ yields the same result. ∎
Appendix F Bounding Total Variation Distance
As $|\Delta|\to n$, we prove that the total variation (TV) distance between the empirical distribution of $z_{\Delta}$ and $z_{[n]}$ goes to 00. Here, recall that $z_{i}\in\mathcal{Z}=\mathcal{S}_{l}\times\mathcal{A}_{l}$, and $z_{\Delta}=\{z_{i}:i\in\Delta\}$ for $\Delta\in{[n]\choose k}$. Before bounding the total variation distance between $F_{z_{\Delta}}$ and $F_{z_{\Delta^{\prime}}}$, we first introduce Lemma C.5 of Anand and Qu [2024] which can be viewed as a generalization of the Dvoretzky-Kiefer-Wolfowitz concentration inequality, for sampling without replacement. We first make an important remark.
Remark F.1.
@@ Appendix F Bounding Total Variation Distance
For reference, we present the multidimensional Dvoretzky-Kiefer-Wolfowitz (DKW) inequality (Dvoretzky et al. [1956], Massart [1990], Naaman [2021]) which bounds the difference between an empirical distribution function for a set $B_{\Delta}$ and $B_{[n]}$ when each element of $\Delta$ for $|\Delta|=k$ is sampled uniformly at random from $[n]$ with replacement.
Theorem F.2 (Multi-dimensional Dvoretzky-Kiefer-Wolfowitz (DFW) inequality [Dvoretzky et al., 1956, Naaman, 2021]).
@@
Lemma C.5 of Anand and Qu [2024] generalizes the DKW inequality for sampling without replacement:
Lemma F.3 (Sampling without replacement analogue of the DKW inequality, Lemma C.5 in Anand and Qu [2024]).
@@
Lemma F.4.
With probability atleast $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$,
$\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})\leq\epsilon$
@@
Recall that $z:=(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$. From Lemma F.3, substituting $\mathcal{B}_{l}=\mathcal{Z}_{l}$ yields:
@@
Lemma F.5 (Lemma B.8.1 of Ghosal and van der Vaart [2017]. Subadditivity of TV distance for Product Distributions).
@@
Lemma F.6 (KL-divergence decays too slowly).
@@
From the Bretagnolle-Huber inequality Tsybakov [2008] we have that
@@
In the third line, we note that $\sum_{z\in\mathcal{Z}_{l}}\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}=|\Delta|$ since each local agent contained in $\Delta$ must have some state/action pair contained in $\mathcal{Z}_{l}$. In the last line, we note that $\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\leq\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}$, For all $z\in\mathcal{Z}_{l}$, and thus the summation of logarithmic terms in the third line is negative.
Finally, using this bound in the Bretagnolle-Huber inequality yields the lemma.∎
Corollary F.7.
From Lemma F.6, setting $\Delta=[n]$ also recovers $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})=0$.
Theorem F.8.
@@
$\left|\hat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\hat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
@@
From combining the total variation distance bound in Lemma F.4 and the Lipschitz continuity bound in Theorem E.3 with $\sum_{t=0}^{T}\gamma^{T}\leq\frac{1}{1-\gamma}$ for $\gamma\in(0,1)$, we have:
$\Pr\left[|\hat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\hat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})|\leq\frac{2\epsilon}{1-\gamma}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}\right]\geq 1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$
Then, reparameterizing $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$ into $1-\delta$ to get $\epsilon=\sqrt{\frac{n-k+1}{8kn}\ln\left(\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}\right)}$ gives that with probability at least $1-\delta$,
$\left|\hat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\hat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty},$
@@
Appendix G Using the Performance Difference Lemma
In general, convergence analysis requires the guarantee that a stationary optimal policy exists. Fortunately, when working with the empirical distribution function, the existence of a stationary optimal policy is guaranteed when the state/action spaces are finite or countably infinite. However, lifting the knowledge of states onto the continuous empirical distribution function space, and designing a policy on the lifted space is still analytically challenging. To circumvent this, Gu et al. [2021] creates lifted $\epsilon$-nets and does kernel regression to obtain convergence guarantees. Moreover, our result has a similar flavor to MDPs with dynamic exogenous inputs from learning theory, [Dietterich et al., 2018, Foster et al., 2022, Anand and Qu, 2024], wherein our subsampling algorithm treats each sampled state as an endogenous state.
Here, our analytic approach bears a stark difference, wherein we analyze the sampled structure of the mean-field empirical distribution function, rather than studying the structure of the lifted space. For this, we leverage the classical Performance Difference Lemma, which we restate below for completeness.
Lemma G.1 (Performance Difference Lemma, Kakade and Langford [2002]).